If 38.05 mL of is required to titrate 2.630 g of Na2CO3 according to the following equation, what is the concentration (mol/L) of the HCL solution?
Na2CO3(aq)+2HCl(aq)--->2NaCl(aq)+CO2(g)+H2O(l)
Concentration = _______ M
1
Expert's answer
2020-05-11T14:24:02-0400

1) 106g___________________73g

2,630g_________________X=2,630*73/106=1,811g HCI

2) C(HCI)=1,811*1000/36,5*38,05=1,3 M